6. A company uses activity-based costing to determine the costs of its three products: A, B,...

6.

A company uses activity-based costing to determine the costs of its three products: A, B, and C. The budgeted cost and activity for each of the company's three activity cost pools are shown in the following table:

Budgeted Activity
Activity Cost Pool Budgeted Cost Product A Product B Product C
Activity 1 $ 70,000 6,000 9,000 20,000
Activity 2 $ 45,000 7,000 15,000 8,000
Activity 3 $ 82,000 2,500 1,000 1,625


What are the activity rates for the three activities under activity-based costing?

Group of answer choices

(1) $2.00; (2) $3.00; (3) $3.50.

(1) $3.50; (2) $1.50; (3) $32,80.

(1) $3.50; (2) $3.00; (3) $16.00.

(1) $2.00; (2) $1.50; (3) $32.80.

(1) $2.00; (2) $1.50; (3) $16.00.
Answer #1

Activity rate=Respective cost/Respective activity

Activity rate for:

Activity 1=70,000/(6000+9000+20,000)=$2

Activity 2=45000/(7000+15000+8000)=$1.5

Activity 3=82000/(2500+1000+1625)=$16

Hence the correct option is:

(1) $2.00; (2) $1.50; (3) $16.00.